Operations Report – March 3rd

Crew 223 Operations Report 03/03/2020

SOL: 3

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Monday generator:
Off at 4:30am
On at 9:10pm SOC was 83%

Tuesday generator:
Off at 5:00am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 80%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149678.53 units

Water (static tank) – 350 gallons

Static to Loft Pump used – yes

Water in GreenHab – 201.4 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes
Hours: 149.4
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: yes
Hours: 133.6
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: no
Hours: 139.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Run during the engineer EVA for battery
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: The ATV 1 and 2’s oil light up when starting the ATV and
then turn off.

HabCar used and why, where? Nothing to report.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 4, 8, 9, 10, 2 were used during the
EVA. We had troubles with suit number 10 : the fans stopped after 26
minutes. We think this is because the suit was not well plugged and thus
was not charged. We are checking this and will continue tomorrow.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Following Scott Davis’ recommendations, we
tried to fix suit 7. We checked most of the electrical connection points
and it doesn’t seem to be a loose electrical connection. We tend for a
problem with the suit’s battery (but we need to confirm this guess
tomorrow), and we will also remove the fans to make sure they are not
obstructed.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Do you know were is
the key of the upper black storage box in the RAM (I would like to open
the drawer containing the wire strippers) ? I only found the one opening
the lower black storage box.